# Simplifying Debiased Inference via Automatic Differentiation and Probabilistic Programming

### v1

@@ G.2.1   Conditional mean
G.2.2   Multifold conditional mean
As in Appendix D.2.2, $\theta(P,u)(\cdot)=E_{P^{r}}[u(Z_{[r]})\,|\,X_{[r]}=\cdot\,]$ and $\dot{\theta}_{P,u}^{*}(w)=(\dot{\nu}_{P,u}^{*}(w),\dot{\zeta}_{P,u}^{*}(w))$, where $\dot{\zeta}_{P,u}^{*}(w)(z_{[r]})=w(x_{[r]})$ and
@@ G.2.2   Multifold conditional mean
In the forward routine, $\theta(P,u)$ can be estimated by regressing $u(Z_{i[r]})$ against $(X_{i[r]})$, with $i[r]:=(i(1),i(2),\ldots,i(r))$ varying over the set $\mathscr{I}_{r}^{(1)}$ of $n^{(1)}!/(n^{(1)}-r)!$ choices of ordered tuples of $r$ unique elements from the first fold $\mathscr{I}^{(1)}$; to reduce runtime, a random subsample of the indices in $\mathscr{I}_{r}^{(1)}$ may be used while fitting this regression. Just as in Appendix D.2.1, this estimate, $\widehat{h}$, will belong to $\theta(\mathcal{M},u)$ as long as it respects any known bounds on $u$ and $\mathcal{M}$ is suitably large.
@@
G.2.3   Conditional covariance
As in Appendix D.2.3, $\theta(P,u)(x)=\mathrm{cov}_{P}[u_{1}(Z),u_{2}(Z)\,|\,X=x]$ and $\dot{\theta}_{P,u}^{*}(w)=(\dot{\nu}_{P,u}^{*}(w),\dot{\zeta}_{P,u}^{*}(w))$, where
@@ G.2.3   Conditional covariance
G.2.4   Conditional variance
When $\theta(P,u)(x)=\mathrm{Var}_{P}[u(Z)\,|\,X=x]$ as in Appendix D.2.4, the nuisance estimation routines for the conditional covariance can be used with $u_{1}=u_{2}=u$ (see Appendix G.2.3).
G.2.5   Kernel embedding
@@
G.3   Maps that only depend on their distribution-valued argument
G.3.1   Root-density
@@
G.3.2   Conditional density
As in Appendix D.3.3, $\theta(P,u)(z)=p_{Y\,|\,X}(z)$ and $\dot{\nu}_{P}^{*}(w)(z)=w(z)p_{Y|X}(z)-E_{P}[w(Z)p_{Y|X}(Z)\,|\,X=x]$. In the forward routine, the estimate $\widehat{h}$ of $p_{Y|X}$ can be obtained using kernel density estimation, if $\lambda$ is a Lebesgue measure, or nonparametric multinomial logistic regression, if $\lambda$ is a counting measure. In either case, the backward routine sets $\widehat{\vartheta}=(\widehat{\vartheta}_{0},0)$ with $\widehat{\vartheta}_{0}(z)=w(z)\widehat{h}(z)-\int w(x,\tilde{y})\widehat{h}(x,\tilde{y})\lambda(d\tilde{y})$. The estimates $\widehat{\vartheta}$ and $\widehat{h}$ are plug-in estimators based on any distribution $\widehat{P}$ whose conditional density of $Y|X$ relative to $\lambda$ is $\widehat{h}$, and so the compatibility condition will hold if the nonparametric model $\mathcal{M}$ is large enough to contain such a distribution.
G.3.3   Dose-response function
As in Appendix D.3.4, $\theta(P,u)(a)=\int\mu_{P}(a,x)P_{X}(dx)$ and
@@ G.3.4   Counterfactual density
G.4   Maps that only depend on their Hilbert-valued argument
G.4.1   Overview
We now consider primitives that only depend on their Hilbert-valued input in the sense described in Appendix D.4.1, so that there is a Hadamard differentiable map $\zeta:\mathcal{U}\rightarrow\mathcal{W}$ such that $\theta(P^{\prime},u^{\prime})=\zeta(u^{\prime})$ for all $(P^{\prime},u^{\prime})\in\mathcal{M}\times\mathcal{U}$.
@@ G.4.1   Overview
G.4.2   Pointwise operations
The adjoint $\dot{\zeta}_{u}^{*}(w)$ for the pointwise operation primitive from Appendix D.4.5 is invariant to the choice of $\lambda$ indexing the ambient Hilbert spaces $\mathcal{T}=L^{2}(\lambda)^{\oplus d}$ and $\mathcal{W}=L^{2}(\lambda)$. Therefore, if $\lambda=P$, then the backward nuisance routine can set $\widehat{\vartheta}$ equal to its true value, $\dot{\theta}_{P,u}^{*}(w)=(0,\dot{\zeta}_{u}^{*}(w))$.
G.4.3   Bounded affine map
We present a backward nuisance routine for the bounded affine maps $\zeta(\cdot)=\kappa(\cdot)+c$ from Appendix D.4.6 in cases where the form of the adjoint $\kappa^{*}$ is unknown. This routine is based on a Riesz loss (Chernozhukov et al., 2022c). Recalling that $\dot{\zeta}_{u}=\kappa$ and $\dot{\zeta}_{u}^{*}=\kappa^{*}$, the key observation used is that $\dot{\zeta}_{u}^{*}(w)=\kappa^{*}(w)$ is the unique solution to
@@ G.4.3   Bounded affine map
$\widehat{\Phi}_{u,w}(t):=\frac{1}{n_{\mathscr{I}}}\sum_{i\in\mathscr{I}}t(Z_{i})-2\frac{1}{n_{\mathscr{I}}}\sum_{i=1}^{n}\kappa(t)(Z_{i})\,w(X_{i}).$
@@
Since the primitives in Appendices D.4.7-D.4.11 are all special cases of bounded affine maps, the backward nuisance routine from this appendix can be used for all of those primitives.

### v3

@@ F.2.1   Conditional mean
F.2.2   Multifold conditional mean
As in Appendix C.2.2, $\theta(P,u)(\cdot)=E_{P^{r}}[u(Z_{[r]})\,|\,X_{[r]}=\cdot\,]$ and $\dot{\theta}_{P,u}^{*}(w)=(\dot{\nu}_{P,u}^{*}(w),\dot{\zeta}_{P,u}^{*}(w))$, where $\dot{\zeta}_{P,u}^{*}(w)(z_{[r]})=w(x_{[r]})$ and
@@ F.2.2   Multifold conditional mean
In the forward routine, $\theta(P,u)$ can be estimated by regressing $u(Z_{i[r]})$ against $(X_{i[r]})$, with $i[r]:=(i(1),i(2),\ldots,i(r))$ varying over the set $\mathscr{I}_{r}^{(1)}$ of $n^{(1)}!/(n^{(1)}-r)!$ choices of ordered tuples of $r$ unique elements from the first fold $\mathscr{I}^{(1)}$; to reduce runtime, a random subsample of the indices in $\mathscr{I}_{r}^{(1)}$ may be used while fitting this regression. Just as in Appendix C.2.1, this estimate, $\widehat{h}$, will belong to $\theta(\mathcal{M},u)$ as long as it respects any known bounds on $u$ and $\mathcal{M}$ is suitably large.
@@
F.2.3   Conditional covariance
As in Appendix C.2.3, $\theta(P,u)(x)=\mathrm{cov}_{P}[u_{1}(Z),u_{2}(Z)\,|\,X=x]$ and $\dot{\theta}_{P,u}^{*}(w)=(\dot{\nu}_{P,u}^{*}(w),\dot{\zeta}_{P,u}^{*}(w))$, where
@@ F.2.3   Conditional covariance
F.2.4   Conditional variance
When $\theta(P,u)(x)=\mathrm{Var}_{P}[u(Z)\,|\,X=x]$ as in Appendix C.2.4, the nuisance estimation routines for the conditional covariance can be used with $u_{1}=u_{2}=u$ (see Appendix F.2.3).
F.2.5   Kernel embedding
@@
F.3   Maps that only depend on their distribution-valued argument
F.3.1   Root-density
@@
F.3.2   Conditional density
As in Appendix C.3.3, $\theta(P,u)(z)=p_{Y\,|\,X}(z)$ and $\dot{\nu}_{P}^{*}(w)(z)=w(z)p_{Y|X}(z)-E_{P}[w(Z)p_{Y|X}(Z)\,|\,X=x]$. In the forward routine, the estimate $\widehat{h}$ of $p_{Y|X}$ can be obtained using kernel density estimation, if $\lambda$ is a Lebesgue measure, or nonparametric multinomial logistic regression, if $\lambda$ is a counting measure. In either case, the backward routine sets $\widehat{\vartheta}=(\widehat{\vartheta}_{0},0)$ with $\widehat{\vartheta}_{0}(z)=w(z)\widehat{h}(z)-\int w(x,\tilde{y})\widehat{h}(x,\tilde{y})\lambda(d\tilde{y})$. The estimates $\widehat{\vartheta}$ and $\widehat{h}$ are plug-in estimators based on any distribution $\widehat{P}$ whose conditional density of $Y|X$ relative to $\lambda$ is $\widehat{h}$, and so the compatibility condition will hold if the nonparametric model $\mathcal{M}$ is large enough to contain such a distribution.
F.3.3   Dose-response function
As in Appendix C.3.4, $\theta(P,u)(a)=\int\mu_{P}(a,x)P_{X}(dx)$ and
@@ F.3.4   Counterfactual density
F.4   Maps that only depend on their Hilbert-valued argument
F.4.1   Overview
We now consider primitives that only depend on their Hilbert-valued input in the sense described in Appendix C.4.1, so that there is a Hadamard differentiable map $\zeta:\mathcal{U}\rightarrow\mathcal{W}$ such that $\theta(P^{\prime},u^{\prime})=\zeta(u^{\prime})$ for all $(P^{\prime},u^{\prime})\in\mathcal{M}\times\mathcal{U}$.
@@ F.4.1   Overview
F.4.2   Pointwise operations
The adjoint $\dot{\zeta}_{u}^{*}(w)$ for the pointwise operation primitive from Appendix C.4.5 is invariant to the choice of $\lambda$ indexing the ambient Hilbert spaces $\mathcal{T}=L^{2}(\lambda)^{\oplus d}$ and $\mathcal{W}=L^{2}(\lambda)$. Therefore, if $\lambda=P$, then the backward nuisance routine can set $\widehat{\vartheta}$ equal to its true value, $\dot{\theta}_{P,u}^{*}(w)=(0,\dot{\zeta}_{u}^{*}(w))$.
F.4.3   Bounded affine map
We present a backward nuisance routine for the bounded affine maps $\zeta(\cdot)=\kappa(\cdot)+c$ from Appendix C.4.6 in cases where the form of the adjoint $\kappa^{*}$ is unknown. This routine is based on a Riesz loss (Chernozhukov et al., 2022c). Recalling that $\dot{\zeta}_{u}=\kappa$ and $\dot{\zeta}_{u}^{*}=\kappa^{*}$, the key observation used is that $\dot{\zeta}_{u}^{*}(w)=\kappa^{*}(w)$ is the unique solution to
@@ F.4.3   Bounded affine map
$\widehat{\Phi}_{u,w}(t):=\frac{1}{n_{\mathscr{I}}}\sum_{i\in\mathscr{I}}t(Z_{i})^{2}-2\frac{1}{n_{\mathscr{I}}}\sum_{i=1}^{n}\kappa(t)(Z_{i})\,w(X_{i}).$
@@
Since the primitives in Appendices C.4.7-C.4.11 are all special cases of bounded affine maps, the backward nuisance routine from this appendix can be used for all of those primitives.